# Evaluating radiation impact on transmon qubits in above and underground facilities

### v1

@@ V.1 Device Fabrication
The devices are fabricated on a 432 m thick -plane HEMEX grade double-polished sapphire substrate with 2-inch diameter. The substrate preparation, base-layer (Nb) deposition and Josephson junction evaporation are similar to the fabrication methods followed in Ref. 43. In order to prevent oxidation of the Niobium, a capping layer of thickness nm is deposited using the same sputtering tool without breaking the vacuum. Gold and Tantalum are used as capping layers for devices Q1 and Q2-Q4 respectively (on two different chips). Device Q1 is first measured at FNAL and then sent to LNGS. The other qubit samples (Q2-Q4) are measured at FNAL. In Table S1, we list the measured parameters of the devices and Fig. S1 shows the relative position of those.
@@
Qubit pulse length | 0.150 | 0.091 | 0.124 | 0.160 | s |
Qubit average T1 | 84 | 141 | 131 | 214 | s |
Readout pulse length | 4.5 | 3.8 | 4.0 | 8.0 | s |
Waiting period | 5.0 | 10.0 | 5.0 | 5.0 | s |
Cooldown period | 50.0 | 70.0 | 70.0 | 10.0 | s |
One iteration period | 64.550 | 87.929 | 84.324 | 31.660 | s |
@@ V.2 Measurement Setup
Our measurement setups at the two locations are shown in Fig. S2. The transmon qubit is measured in the hangar geometry, where a CPW readout resonator is coupled to a transmission line. The pulses for both qubit control and readout are sent through a common input line. The input line is attenuated by 60-70 dB distributed between the 4-K and mixing chamber stages. Copper-body 20 dB attenuators are used at the base plate plate for better thermalization. At LNGS, the input line is also filtered through a 8 GHz reflective low-pass filter and a strong infrared filter (QMC-CRYOIRF-002) to reduce high-frequency noise. The superconducting output line is protected from the incoming noise through a 4-8 GHz triple junctions isolator, an 8 GHz low-pass filter and a weak infrared filter (QMC-CRYOIRF-001). At FNAL, both the input and output lines are filtered through 11 GHz reflective low-pass filters. The superconducting output line also has two 4-8 GHz isolators. The device boxes at both locations are protected from stray magnetic fields using cryoperm shields. The output signal is first amplified through a commercial HEMT amplifier at the 4-K plate, followed by a room-temperature amplifier outside the fridge before digitization. The qubit control and readout pulses are generated digitally using a RFSoC ZCU 216 evaluation board and the QICK [49] software.
@@ V.3 Simulation of Interaction Rates
A two layers Permimphy magnetic shielding (Outer layer: 55.6 cm (OD) 91.3 cm, 1.5 mm-thick; Inner layer: 51 cm (ID) 89 cm, 1.5 mm-thick);
@@
A 48.1 cm (OD) 119.5 cm simplified version of the cryostat with internal vessels;
@@ A CryoPerm® magnetic shield below the mixing chamber;
Figure S4 presents the results of the Monte Carlo simulation, showing the rate of events as a function of the energy deposited in the qubit substrate for both FNAL and INFN-LNGS facilities. The event rates for each radiation source are reported in Table 1. Notably, the overall rate is expected to be approximately an order of magnitude lower at INFN-LNGS (4 @INFN-lNGS vs. 57 @FNAL), where the dominant contribution is from the PCBs.
To investigate the rate in the presence of the Thorium radioactive sources used at INFN-LNGS, we generate radioactive decays from a point-like 232Th radioactive source positioned between the external magnetic shield and the copper bars, approximately at the same height of the chip. For this simulation, we consider only gamma emitters in the chain (228Ra, 228Ac, 212Pb, 212Bi and 208Tl), since and particles cannot reach the chip as they are shielded by the cryostat. As in the previous case, we determine the rate of interactions in the chip by scaling the number of recorded events to the activities reported in Table S3, which correspond to the Thorium source activities employed in the measurements. The rate of events in the chip, plotted as a function of the deposited energy for the various activities, is illustrated in Fig. S3.
@@ V.4 Qubit State Discrimination
The I/Q values (in-phase and quadrature-phase signals) obtained for each measurement depend on the state of the qubit. When a large counts of measurements are plotted on the I-Q plane, we usually observe two blobs corresponding to and . However, we sometimes observe signals corresponding the transmon being in the second excited state as shown in Fig. S6 and we discard those measurements in the analysis.
The blobs corresponding to and are rotated in order to have the centers at the same value of Q and projected onto the I axis. In this way it is possible to discriminate between the two states by putting a threshold value only on one (I) quadrature.
The populations of the states have been estimated by doing a simultaneous fit on the two quadratures with the RooFit toolkit. The fit function used was a sum of 2D gaussians, one for each blob we were observing. The data was divided into traces of approx. 1 minute duration, and for each of them the simultaneous fit was performed. In this way we were able to identify periods in which the qubit was showing a high population of the state and reject them.
V.5 Data Analysis
In our analysis we selected events where the probability of observing a certain number of zeros and ones, calculated according to the Binomial distribution, was below a predefined threshold. Specifically, if the probability of observing such a series of zeros and ones fell below 1%, corresponding to an absolute logarithm () greater than 4.6, the software trigger was fired, indicating the presence of a potential signal. Once triggered, these data wedre recorded for further analysis.
By selecting pulses with , we observe a trigger rate of 5–6 events/second in the runs performed with the radioactive sources, significantly higher than the expected rate from radioactivity (0.1–0.4 events per second depending on the source activity), suggesting that the majority of triggers are due to noise. To address this, we computed additional variables to refine our data selection. Specifically, we calculated:
@@ V.5 Data Analysis
The value and RMS of the baseline before the pulse;
@@
The maximum value of ;
3.
The pulse length, defined as the number of points above threshold;
4.
The position of the maximum.
The baseline and its RMS were computed using the first 100 points preceding the trigger. We then removed pulses occurring on the tail of a preceding pulse or those with a “noisy” baseline (see Fig. 4) by requiring a baseline value lower or equal to its mean value + 3. Additionally, we ensured that the position of the maximum was close to the trigger to avoid re-triggering of the same pulse. Since radioactivity typically results in long pulses (at least 1 ms), we further refined our selection by choosing pulses with length greater than 50 points, corresponding to approximately 3.5 ms under our data-taking conditions.
Finally, we conducted a study to assess the suitability of the window size used for computing the Binomial probability. Our initial choice of 50 points stemmed from the analysis of typical events detected in calibration runs. By averaging tens of these events, we determined that the qubit returned to equilibrium with a characteristic time constant () of 0.78 ms. To ensure complete recovery of the qubit, we adopted a window size corresponding to 5, or approximately 4 ms, which translates to 56 points given our sampling frequency of about 0.07 ms-1. We subsequently performed the analysis using window sizes of 20, 30, 70, and 90 points. Our findings confirmed that, for all these window sizes, the detection efficiency could not reach values above 70. Indeed, when the window size is too small, only a fraction of the signal can be captured within the window. Conversely, if the window size exceeds the signal length, the number of zeros associated with random qubit de-excitation becomes significant, thus deteriorating the signal-to-noise ratio.

### v2

@@ V.1 Device Fabrication
The devices are fabricated on a 432 m thick -plane HEMEX grade double-polished sapphire substrate with 2-inch diameter. The substrate preparation, base-layer (Nb) deposition and Josephson junction evaporation are similar to the fabrication methods followed in Ref. 46. In order to prevent oxidation of the Niobium, a nm thick capping layer is deposited using the same sputtering tool without breaking the vacuum. Gold and Tantalum are used as capping layers for devices Q1 and Q2-Q4 respectively (on two different chips). Device Q1 is first measured at FNAL and then sent to LNGS. The other qubit samples (Q2-Q4) are measured at FNAL. In Table S1, we list the measured parameters of the devices and Fig. S1 shows the relative position of those.
@@
Qubit pulse length | 0.195 | 0.091 | 0.124 | 0.160 | s |
Readout pulse length | 6.0 | 3.8 | 4.0 | 8.0 | s |
@@ V.2 Measurement Setup
Our measurement setups at the two locations are shown in Fig. S2. The transmon qubit is measured in the hangar geometry, where a CPW readout resonator is coupled to a transmission line. The pulses for both qubit control and readout are sent through a common input line. The input line is attenuated by 60-70 dB distributed between the 4-K and mixing chamber stages. Copper-body 20 dB attenuators are used at the base plate for better thermalization. At LNGS, the input line is also filtered through a 8 GHz reflective low-pass filter and a strong infrared filter (QMC-CRYOIRF-002) to reduce high-frequency noise. The superconducting output line is protected from the incoming noise through a 4-8 GHz triple junctions isolator, an 8 GHz low-pass filter and a weak infrared filter (QMC-CRYOIRF-001). At FNAL, both the input and output lines are filtered through 11 GHz reflective low-pass filters. The superconducting output line also has two 4-8 GHz isolators. The device boxes at both locations are protected from stray magnetic fields using cryoperm shields. The output signal is first amplified through a commercial HEMT amplifier at the 4-K plate, followed by a room-temperature amplifier outside the fridge before digitization. The qubit control and readout pulses are generated digitally using a RFSoC ZCU 216 evaluation board and the QICK [50] software.
@@ V.3 Simulation of Interaction Rates
A two layers Permimphy magnetic shielding (Outer layer: 55.6 cm (OD) 91.3 cm (height), 1.5 mm-thick; Inner layer: 51 cm (ID) 89 cm (height), 1.5 mm-thick);
@@
A 48.1 cm (OD) 119.5 cm (height) simplified version of the cryostat with internal vessels;
@@ A CryoPerm® magnetic shield below the mixing chamber;
Figure S4 presents the results of the Monte Carlo simulation, showing the rate of events as a function of the energy deposited in the qubit substrate for both FNAL and INFN-LNGS facilities. The event rates for each radiation source are reported in Table 1. Notably, the overall rate is expected to be approximately an order of magnitude lower at INFN-LNGS (4 event/sec vs. 57 event/sec @FNAL), where the dominant contribution is from the PCBs.
To investigate the rate in the presence of the Thorium radioactive sources used at INFN-LNGS, we generate radioactive decays from a point-like 232Th radioactive source positioned between the external magnetic shield and the copper bars, approximately at the same height of the chip. For this simulation, we consider only gamma emitters in the chain (228Ra, 228Ac, 212Pb, 212Bi and 208Tl), since and particles cannot reach the chip as they are shielded by the cryostat. As in the previous case, we determine the rate of interactions in the chip by scaling the number of recorded events to the activities reported in Table S3, which correspond to the Thorium source activities employed in the measurements (and previously characterized using High Purity Germanium spectroscopy). The rate of events in the chip, plotted as a function of the deposited energy for the various activities, is illustrated in Fig. S3.
@@ V.4 Qubit State Discrimination
The I/Q values (in-phase and quadrature-phase signals) obtained from each measurement reflect the state of the qubit. When a large number of measurements are plotted on the I-Q plane, two clusters, corresponding to the states (ground) and (excited), are observed. In our data we observe also signals corresponding to the qubit being in the second excited state and, occasionally, in the third excited state . To ensure accuracy we discard measurements where the populations of and are too high.
To identify and exclude periods where the qubit exhibits high populations of the and states, we fit the clouds with 2D Gaussian functions, one for each cluster observed. In our experiment the data are divided into traces of measurements. The fit is performed on each trace.
To fit the clouds we use the RooFit toolkit [57]. The fit is performed by projecting the clouds along the two quadratures and simultaneously fitting both projections. The fit function for each projection is given by , where is a Gaussian centered in and with standard deviation . Each Gaussian is normalized such that corresponds to the number of measurements associated with that state. The population of a state is given then by . Fig. S6 shows the clouds produced by a single trace and the result of the simultaneous fit used to estimate state populations.
The clouds corresponding to and are orientated such that their centers are aligned along the same Q value. In this way it is possible to discriminate between the two states by setting a threshold only on the I axis, simplifying state identification.
V.5 “Effective” calculation
To calculate the “effective” relaxation time over a trace of points, we perform the following steps:
@@
We count the number of times the qubit is in the excited state () OVER the total number of measurements ;
@@
We calculate as .
For each dataset, this process is repeated for each segment of points to monitor changes over time, and then we calculate the average , which is used to estimate the probability of the qubit being in state . This method ensures accurate determination of the qubit’s relaxation time by averaging over a large number of measurements.
V.6 Threshold values
To distinguish events caused by ionizing radiation from those due to qubit spontaneous decay, we employed a consistent analysis approach across all datasets collected at both LNGS and FNAL. Our method involves calculating the probability of observing more than a certain number of zeros within a 60-point signal window using the CDF of the Binomial distribution. This allows us to establish dataset-specific thresholds aimed at minimizing the contribution from qubit spontaneous decay while potentially maximizing sensitivity to radiation-induced events. Specifically, these thresholds are set to target spontaneous qubit decay rates approximately an order of magnitude lower than those expected from ambient radiation sources. Additionally, to mitigate noise-induced triggers, a separate threshold is applied to the pre-trigger region, ensuring robustness in our event selection criteria.
Table S4 summarizes the specific thresholds applied to each dataset, alongside the expected rates from qubit spontaneous decay (calculated using these thresholds), the expected rates from ionizing radiation, and the measured rates.
Datasets | Noise | Signal | Qubit Spontaneous | Ionizing Radiation | Measured |
Threshold | Threshold | Decay Rate | Rate | Rate | |
[# of zeros] | [# of zeros] | [events/sec] | [events/sec] | [events/sec] | |
Shielded chip | 15 | 23 | 0.0001 | 0.0040 0.0006 | 0.003 0.001 |
44 kBq Th Source | 14 | 24 | 0.001 | 0.12 0.01 | 0.032 0.002 |
76 kBq Th Source | 14 | 23 | 0.002 | 0.20 0.02 | 0.036 0.004 |
125 kBq Th Source | 15 | 24 | 0.004 | 0.34 0.03 | 0.052 0.006 |
161 kBq Th Source | 15 | 24 | 0.002 | 0.43 0.04 | 0.056 0.004 |
FNAL Q2 | 10 | 18 | 0.002 | 0.057 0.003 | 0.003 0.001 |
FNAL Q3 | 14 | 23 | 0.002 | 0.057 0.003 | 0.005 0.001 |
FNAL Q4 | 10 | 18 | 0.001 | 0.057 0.003 | 0.025 0.004 |